OpenStudy (anonymous):

Gregor has $240 in savings. He plans to add $35 to his savings each month. How much will he have in savings after 6 months? A. $275 B. $281 C. $415 D. $450

OpenStudy (anonymous):

6*35=210 + With his saves makes it: 450. Answer is D

OpenStudy (anonymous):

he has 240, and he adds 35*6 additionally, in a 6 month period. \[240+35(6)=240+210=450~~~~~~~~~~~~~~~~~~~~D~~~~~~☻\]

OpenStudy (anonymous):

hes right dude he get the medal ok sry who answers frist